a) What mass of carbon dioxide is produced from the complete combustion of 3.90×10−3 g of methane? Express your answer with the appropriate units.

b) What mass of water is produced from the complete combustion of 3.90×10−3 g of methane? Express your answer with the appropriate units.

c) What mass of oxygen is needed for the complete combustion of 3.90×10−3 g of methane? Express your answer with the appropriate units.

d) The fuel used in many disposable lighters is liquid butane, C4H10. How many carbon atoms are in 3.00 g of butane? Express your answer numerically in atoms.
Answer #1

a) What mass of carbon dioxide is produced from the complete combustion of 3.90 × 10−3 g of methane? Express your answer with the appropriate units.

CH4 + 2O2 ---> CO2 + 2H2O


Molar mass of CH4 = 12 + 4 x 1 = 16 g / mole
Molar mass of CO2 = 12 + 2 * 16 = 44 g / mole

16 g of CH4 produces 44 g of CO2.
3.90 x 10-3 g of CH4 produces = ( 44 x 3.90 x 10-3) / 16
= 1.07 x 10-2 g of CO2

b) What mass of water is produced from the complete combustion of 3.90×10−3 g of methane? Express your answer with the appropriate units.

CH4 + 2O2 ---> CO2 + 2H2O


Molar mass of CH4 = 12 + 4 x 1 = 16 g / mole
Molar mass of H2O is = 2 x 1 + 16 = 18 g / mole

16 g of CH4 produces = 2 x 18 = 36 g of H2O

3.90 x 10-3 g of CH4 produces = ( 3.90 x 10-3 x 36) / 16
= 8.78 x 10-3 g of H2O

c) What mass of oxygen is needed for the complete combustion of 3.90×10−3 g of methane? Express your answer with the appropriate units.​​​​​​​


CH4 + 2O2 ---> CO2 + 2H2O


Molar mass of CH4 = 12 + 4 x 1 = 16 g / mole
Molar mass of O2 is = 2 x 16 = 32 g / mole

16 g of CH4 reacts with 2 x 32 g = 64 g of O2
3.90 x 10-3 g of CH4 reacts = ( 64 x 3.90 x 10-3) / 16
= 1.56 x 10-2 g of O2

d) The fuel used in many disposable lighters is liquid butane, C4H10. How many carbon atoms are in 3.00 g of butane? Express your answer numerically in atoms.​​​​​​​

1 mole of butane has 4 mole of carbon atom.

58 grams of C4H10 contains = 4 x 6.023 x 1023 atoms of carbon

3.0 grams of C4H10 contains = (3 x 4 x 6.023 x 1023) / 58 atoms of carbon

= 1.25 x 1023 atoms of carbon